How many onces of 5% Hydrochloric acid, 20% hydro. acid, and water must be combined to get 10 oz of solution that is 8.5% hydro. acid if the amount of water used must equal the total amount of the other two solutions?
Answer by ankor@dixie-net.com(12704)

How many ounces of 5% Hydro-chloric acid, 20% hydro acid, and water must be
combined to get 10 oz of solution that is 8.5% hydro acid if the amount of
water used must equal the total amount of the other two solutions?
:
That means that half the final solution has to be water: 5 oz
And half has to be a mixture of the two acids.
When the two are combined the percent will be halved
Calculate the amt required to make 5 oz of 17% solution
:
Let x = amt of 5% acid
then
(5-x) = amt of 20% acid
:
The equation:
.05x + .2(5-x) = .17(5)
.05x + 1 - .2x = .85
.05x - .20x = .85 - 1
-.15x = -.15
:
x = 1 oz of 5% solution
then
4 - 1 = 4 oz of 20% solution
:
When the 5 oz mixture is mixed with 5 oz of water:
.17(5) = .085(5 + 5)
.85 = .085(10)
.85 = .85